OpenStudy (anonymous):

please help! I need to pass algebra to graduate! widget wonders produces widgets. They have found that the cost, c(x), of making x widgets is a quadratic function in terms of x. the company also discovered that it cost $16 to produce 4 widgets, ans $48 to produce 10 widgets. find the total cost of producing 8 widgets.

OpenStudy (curry):

So 16 = a(4)^2 - b4 -c and 48 = a(10)^2 -b(10) - c

OpenStudy (curry):

Correction, those minus signs should be positive. so solve for those two equations, and the Cs cancel immediately. so your'e left with 16 =16a+4b. and 48 = 100a+10b.

OpenStudy (curry):

you can figure out a and b frmo there, and then you can find c once you have a and b. Lastly, when you have the three constants, (a,b,c), you can just plug in 8 for x into the quadratic eqauation. let me know if you need more help.

OpenStudy (anonymous):

yeah Im so lost I don't even know what this means :(

OpenStudy (anonymous):

@Curry

OpenStudy (anonymous):

do you still need help?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok so can you tell me what the general form of a quadratic function is?

OpenStudy (anonymous):

if not thats ok, i need to see what you know and what you dont

OpenStudy (anonymous):

I don't

OpenStudy (anonymous):

ok, the general form of a quadratic function is this ax^2 + bx + c a = some number that is not 0 b = some number c = some number for example 2x^2 + 5x + 4

OpenStudy (anonymous):